The 2014 balance sheet of Greystone, Inc., showed current assets of $2332 and current liabilities of $1723. The 2015 balance sheet showed current assets of $2171 and current liabilities of $1234. What was the company’s 2015 change in net working capital, or NWC
Answer #1

net working capital=current assets-current liabilities

net working capital for 2015=(2171-1234)=$937

net working capital for 2014=(2332-1723)=$609

Hence change in net working capital=(937-609)

=$328